# Double star arrangement and the pointed multinet

### v1

@@ Abstract.
Let $\mathcal{A}$ be a hyperplane arrangement in a complex projective space. It is an open question if the degree one cohomology jump loci (with complex coefficients) are determined by the combinatorics of $\mathcal{A}$. By the work of Falk and Yuzvinsky [FY07], all the irreducible components passing through the origin are determined by the multinet structure, which are combinatorially determined. Denham and Suciu introduced the pointed multinet structure to obtain examples of arrangements with translated positive-dimensional components in the degree one cohomology jump loci [DS14]. Suciu asked the question if all translated positive-dimensional components appear in this manner [Suc14A]. In this paper, we show that the double star arrangement introduced by Ishibashi, Sugawara and Yoshinaga [ISY22, Example 3.2] gives a negative answer to this question.
@@ 1.1. Cohomology jump loci
Assume that $X$ is a connected finite CW complex. Let $\mathrm{Char}(X)$ denote the group of $\mathbb{C}$-valued characters $\mathrm{Hom}(\pi_{1}(X),\mathbb{C}^{*})$, which is a commutative affine algebraic group. Each character $\rho\in\mathrm{Char}(X)$ defines a rank one $\mathbb{C}$-local system on $X$, denoted by $L_{\rho}$.
@@
The degree one cohomology jump loci of $X$ are defined as
@@ 1.2. Multinet
We list some known results due to Pereira and Yuzvinsky as follows.
@@ 1.3. Pointed multinet
We show in next section that the double star arrangement introduced by Ishibashi, Sugawara and Yoshinaga in [ISY22, Example 3.2] gives a negative answer to Question 1.10.
Acknowledgments. The first author would like to thank Professor Masahiko Yoshinaga for valuable discussion. Both authors are supported by National Key Research and Development Project SQ2020YFA070080, NSFC grant No. 12001511, the starting grant from University of Science and Technology of China, the Project of Stable Support for Youth Team in Basic Research Field, CAS (YSBR-001), the project “Analysis and Geometry on Bundles” of Ministry of Science and Technology of the People’s Republic of China and Fundamental Research Funds for the Central Universities.
@@ 2. The double star arrangement
(1) $h_{1}-h_{2}=\frac{\sqrt{5}}{c}\cdot(x^{2}+y^{2}-1)^{2}.$
This shows that $h$ has a multiple fiber of multiplicity 2 at the point $1\in\mathbb{C}^{*}$. So $h$ is a small orbifold fibration with exactly one multiple fiber of multiplicity 2, which gives a translated 1-dimensional component of $\mathcal{V}^{1}(X)$. Since the degree 2 polynomial $x^{2}+y^{2}-1$ appearing in the equality (1) can not decompose as a product of two linear polynomial, this particular translated 1-dimensional component can not come from any pointed multinet structure. ∎
@@
According to the computations in [ISY22, Example 3.2] and the fact that $h$ is small, the second cohomology of this Aomoto complex must have $2$-torsion. Then [LL21, Proposition 1.12] shows that the map $h$ may have a multiple fiber with multiplicity dividing by $2$.

### v2

@@ Abstract.
Let $\mathcal{A}$ be a hyperplane arrangement in a complex projective space. It is an open question if the degree one cohomology jump loci (with complex coefficients) are determined by the combinatorics of $\mathcal{A}$. By the work of Falk and Yuzvinsky [FY07] and Marco-Buzunáriz [Mar09], all the irreducible components passing through the origin are determined by the multinet structure, which is combinatorially determined. Denham and Suciu introduced the pointed multinet structure, which is combinatorially determined, to obtain examples of arrangements with translated positive-dimensional components in the degree one cohomology jump loci [DS14]. Suciu asked the question if all translated positive-dimensional components appear in this manner [Suc14A]. In this paper, we show that the double star arrangement introduced by Ishibashi, Sugawara and Yoshinaga [ISY25, Example 3.2] gives a negative answer to this question.
@@ 1.1. Cohomology jump loci
Assume that $X$ is a connected finite CW complex. Let $\mathrm{Char}(X)$ denote the group of $\mathbb{C}^{*}$-valued characters $\mathrm{Hom}(\pi_{1}(X),\mathbb{C}^{*})$, which is a commutative affine algebraic group. Each character $\rho\in\mathrm{Char}(X)$ defines a rank one $\mathbb{C}$-local system on $X$, denoted by $L_{\rho}$.
@@
The degree one cohomology jump loci of $X$ is defined as
@@ 1.2. Multinet
The notion of multinet was simultaneously introduced by Marco-Buzunáriz in [Mar09] by the name of combinatorial pencil. Next we list some known results about multinet due to Pereira and Yuzvinsky as follows.
@@ 1.3. Pointed multinet
In next section, we show that the double star arrangement introduced by Ishibashi, Sugawara and Yoshinaga in [ISY25, Example 3.2] admits a small orbifold map whose multiple fiber is not formed by lines. This example shows one of the difficult points to prove that the translated positive-dimensional component of $\mathcal{V}^{1}(\mathcal{A})$ are combinatorially determined: the multiple fiber of the orbifold map may have irreducible fiber with degree greater than one. In particular, it gives a negative answer to Question 1.10.
Note that Theorem 1.6 implies that for a line arrangement $\mathcal{A}$ a translated positive-dimensional component of $\mathcal{V}^{1}(\mathcal{A})$ may have dimension 1, 2 or 3. The pointed multinet structure can only produce translated 1-dimensional component in $\mathcal{V}^{1}(\mathcal{A})$. The double star arrangement we will studied later has translated 1-dimensional component in $\mathcal{V}^{1}(\mathcal{A})$ not coming from the pointed multinet structure. It is an interesting question if one can find examples of line arrangement $\mathcal{A}$ such that $\mathcal{V}^{1}(\mathcal{A})$ has a translated component with dimension 2 or 3.
Acknowledgments. The first author would like to thank Professor Masahiko Yoshinaga for valuable discussion. Both authors are supported by National Key Research and Development Project SQ2020YFA070080, the Project of Stable Support for Youth Team in Basic Research Field, CAS (YSBR-001), the project “Analysis and Geometry on Bundles” of Ministry of Science and Technology of the People’s Republic of China and Fundamental Research Funds for the Central Universities.
@@ 2. The double star arrangement
$h_{1}-h_{2}=\frac{\sqrt{5}}{c}\cdot(x^{2}+y^{2}-1)^{2}.$
Let $\overline{h_{1}}$ and $\overline{h_{2}}$ be the homogenization of $h_{1}$ and $h_{2}$, respectively. Then we have
(1) $\overline{h_{1}}-\overline{h_{2}}=\frac{\sqrt{5}}{c}z\cdot(x^{2}+y^{2}-z^{2})^{2},$
where $\{z=0\}$ is the line at infinity. This shows that the map $h$ has a multiple fiber of multiplicity 2 at the point $1\in\mathbb{C}^{*}$. So $h$ is a small orbifold fibration with exactly one multiple fiber of multiplicity 2, which gives a translated 1-dimensional component of $\mathcal{V}^{1}(\mathcal{A})\subseteq\mathrm{Char}(X(\mathcal{A}))\cong(\mathbb{C}^{*})^{10}$. This translated 1-dimensional component is of the form $\rho T$, where $\rho=(1,1,1,1,1,-1,-1,-1,-1,-1)$ and $T=(t,t,t,t,t,t^{-1},t^{-1},t^{-1},t^{-1},t^{-1})$. Here we order the affine lines as in Figure 2.
Since the degree 2 polynomial $x^{2}+y^{2}-z^{2}$ appearing in the equality (1) can not decompose as a product of two linear polynomial, this particular translated 1-dimensional component can not come from any pointed multinet structure. ∎
@@
According to the computations in [ISY25, Example 3.2] and the fact that $h$ is small, the second cohomology of this Aomoto complex must have $2$-torsion. Then [LL21, Proposition 1.12] shows that the map $h$ may have a multiple fiber with multiplicity dividing by $2$.